OpenStudy (anonymous):

Find the solution to each system by the addition (elimination) method. Check your answer for exercises 23-26. 5s+9r=6 4s-3t=15

OpenStudy (zzr0ck3r):

s t and r?

OpenStudy (abb0t):

poop on it

OpenStudy (anonymous):

It was need my book

OpenStudy (zzr0ck3r):

say again?

OpenStudy (zzr0ck3r):

this is two equations with three unknowns

OpenStudy (anonymous):

I don't know I just put everything that was in my book on the web site

OpenStudy (zzr0ck3r):

are you sure there is a t and a r?

OpenStudy (anonymous):

sorry I miss it up it is 5s+9t=6 and 4s-3t=15. I'm sorry about that

OpenStudy (zzr0ck3r):

ok this is good:0

OpenStudy (zzr0ck3r):

5s+9t = 6 4s - 3t = 15 we want to multiply one of the equations by something so that when we add it to the other one, one variable drops, multiply the second equation by 3

OpenStudy (zzr0ck3r):

5s + 9t = 6 12s -9t = 45 now add the equations

OpenStudy (anonymous):

and its two liner equations in two variables is to add the two equations

OpenStudy (zzr0ck3r):

17s + 0 = 51 so s = 3 now use either equation to find t

OpenStudy (zzr0ck3r):

5s+9t = 6 plug in s 5(3) + 9t = 6 so t = -1 so s = 3 t = -1 is the only solution

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

not really

OpenStudy (zzr0ck3r):

what part?

OpenStudy (anonymous):

How did you get 3 for s